# RTG-SLAM: Real-time 3D Reconstruction at Scale using Gaussian Splatting

### v1

@@ Abstract.
We propose RTG-SLAM, a real-time 3D reconstruction system with an RGBD camera for large-scale environments using Gaussian splatting. RTG-SLAM features a compact Gaussian representation and a highly efficient on-the-fly Gaussian optimization scheme. We force each Gaussian to be either opaque or nearly transparent, with the opaque ones fitting the surface and dominant colors, and transparent ones fitting residual colors. By rendering depth in a different way from color rendering, we let a single opaque Gaussian well fit a local surface region without the need of multiple overlapping Gaussians, hence largely reducing the memory and computation cost. For on-the-fly Gaussian optimization, we explicitly add Gaussians for three types of pixels per frame: newly observed, with large color errors and with large depth errors. We also categorize all Gaussians into stable and unstable ones, where the stable Gaussians are expected to well fit previously observed RGBD images and otherwise unstable. We only optimize the unstable Gaussians and only render the pixels occupied by unstable Gaussians. In this way, both the number of Gaussians to be optimized and pixels to be rendered are largely reduced, and the optimization can be done in real time. We show real-time reconstructions of a variety of real large scenes. Compared with the state-of-the-art NeRF-based RGBD SLAM, our system achieves comparable high-quality reconstruction but with around twice the speed and half the memory cost, and shows superior performance in the realism of novel view synthesis and camera tracking accuracy.
@@ 1. Introduction
In this paper, we introduce RTG-SLAM, a real-time 3D reconstruction system with an RGBD camera for large-scale environments using Gaussian splatting, featuring a compact Gaussian representation and a highly efficient on-the-fly Gaussian optimization scheme. In our compact Gaussian representation, we force each Gaussian to be either opaque or nearly transparent, with the opaque ones fitting the surface (i.e., depth map) and dominant colors, and transparent ones fitting residual colors. Our intention is to use a single opaque Gaussian to fit a local region of the surface without the need for multiple overlapping Gaussians. However, even for an opaque Gaussian, rendering its depth in the same way as rendering color would produce varying depth values declined from the Gaussian center, making it inaccurate to represent a local area using this Gaussian alone. To this end, we propose to render depth in a different way from color rendering. Following classical point rendering techniques (Zwicker et al., 2001), we treat each opaque Gaussian as an ellipsoid disc on the dominant plane of Gaussian, so that it can well fit a local region or a large flat area by itself. The depth rendering is very convenient under this setting. During color rendering, we already have the sorted Gaussians as well as their opacities for each pixel. By selecting from front to back the first Gaussian whose opacity for the pixel is larger than a given threshold, we consider the ray hits the ellipsoid disc and compute the intersection point using equations of the ray and disc plane. Then the depth for the pixel is equal to the depth of the intersection point. The whole process is differentiable, so Gaussians can be optimized by measuring the differences between the rendered and input depth maps through backpropagation. The compact Gaussian representation can fit the 3D surfaces with much fewer Gaussians, hence largely reducing the memory and computation cost.
@@
We show real-time reconstructions of a variety of real large scenes, including corridor, storeroom, hotel room, home and office, ranging from $43m^{2}\sim 100m^{2}$. All the results are scanned and reconstructed with a Microsoft Azure Kinect in real time (around 16 fps) without any post-processing. Comparisons demonstrate that RTG-SLAM runs at around twice the speed of the state-of-the-art NeRF-based SLAM, with around half the memory cost (e.g., 17.9 fps, 8.8 GB versus 8.65 fps, 17.3 GB (Wang et al., 2023) on the home scene). We also compare our method with the concurrent Gaussian SLAM work SplaTAM (Keetha et al., 2023) (the only one with code published). We also surpass SplaTAM in speed and memory, where SplaTAM runs at 0.31 fps and is out of memory during scanning of the home scene. We also conduct extensive experiments on three widely-used datasets: Replica (Straub et al., 2019), TUM-RGBD (Sturm et al., 2012) and ScanNet++ (Yeshwanth et al., 2023). Compared with the state-of-the-art NeRF SLAM methods, our system achieves comparable high-quality reconstruction, and shows superior performance in time and memory performance, realism of novel view synthesis, and camera tracking accuracy. The code and data are at https://gapszju.github.io/RTG-SLAM.
@@ 2. Related Work
Concurrent with our work, some efforts also explore integrating 3D Gaussians into dense RGBD SLAM. 3D Gaussians (Kerbl et al., 2023) are the state-of-the-art representation of radiance fields that can render high-quality images in real time, but the optimization is conducted offline typically requiring several minutes. To extend Gaussians to online reconstruction, (Yan et al., 2023) proposes an adaptive expansion strategy to add new or delete noisy 3D Gaussian and a coarse-to-fine technique to select reliable Gaussians for tracking. (Yugay et al., 2023) proposes novel strategies for seeding and optimizing Gaussian splats to extend their use to sequential RGBD inputs. SplaTAM (Keetha et al., 2023) tailors an online reconstruction pipeline to use an underlying Gaussian representation and silhouette-guided optimization via differentiable rendering. (Matsuki et al., 2023) unifies the Gaussian representation for accurate, efficient tracking, mapping, and high-quality rendering. (Huang et al., 2023b) introduces a Gaussian-Pyramid-based training method to progressively learn multi-level features and enhance mapping performance. While promising results are demonstrated, it is still difficult in reaching real-time reconstruction at scale. The reported fastest reconstruction speed is 8.34 fps on an NVIDIA RTX 4090 GPU (Matsuki et al., 2023) on the synthetic Replica dataset (Straub et al., 2019). They did not present the complete reconstruction results on real large scenes either. Thanks to our compact Gaussian representation and highly efficient Gaussian optimization strategy, our method can reconstruct real large scenes in real time with low memory cost.
@@ 4.1. Experimental Setup
We implemented our SLAM system on a desktop computer with an intel i9 13900-KF CPU and an Nvidia RTX 4090 GPU. We implemented the mapping and tracking parts in Python using Pytorch framework and wrote custom CUDA kernels for rasterization and back propagation. We used an Azure Kinect as the RGBD camera for real-time scanning. Please refer to the supplementary material for more details.
@@
ScanNet++ is a large-scale dataset that couples together capture of high-quality and commodity-level geometry and color of indoor scenes. Its depth maps are rendered from models reconstructed from laser scanning. Different from other benchmarks, each camera pose in ScanNet++ is very far apart from one another. We also scanned real-world scenes by ourselves to build an Azure dataset, including corridor, storeroom, hotel room, home, office, ranging from $43m^{2}\sim 100m^{2}$.
@@ 4.2. Evaluation of Online Reconstruction
We qualitatively compare the rendering quality for novel view synthesis for all methods. The results are shown in Fig. 7. Note the NeRF-based methods require a depth map to synthesize high-quality images, so we use the reconstructed mesh to render depth maps for them. We also quantitatively compare the novel-view synthesis on the ScanNet++ testing views, where the ground truth depth is used for NeRF-based methods, and the results are reported in the supplementary material. We can see that our method and SplaTAM (2023) clearly produces higher quality images with much fewer artifacts and higher fidelity appearance. We also quantitatively compare the rendering quality with other methods on the training views of Replica, following Point-SLAM (2023) and all concurrent Gaussian SLAM works. Please see the table in the supplementary material. Our method achieves a rendering quality comparable with SplaTAM and Point-SLAM (2023) (which needs the ground-truth depth map as input), and consistently outperforms the other NeRF SLAM methods.
@@ Acknowledgements.
The authors would like to thank the reviewers for their insightful comments. This work is supported by NSF China (No. U23A20311 & 62322209), the XPLORER PRIZE, and the 100 Talents Program of Zhejiang University.

### v2

@@ Abstract.
We present Real-time Gaussian SLAM (RTG-SLAM), a real-time 3D reconstruction system with an RGBD camera for large-scale environments using Gaussian splatting. The system features a compact Gaussian representation and a highly efficient on-the-fly Gaussian optimization scheme. We force each Gaussian to be either opaque or nearly transparent, with the opaque ones fitting the surface and dominant colors, and transparent ones fitting residual colors. By rendering depth in a different way from color rendering, we let a single opaque Gaussian well fit a local surface region without the need of multiple overlapping Gaussians, hence largely reducing the memory and computation cost. For on-the-fly Gaussian optimization, we explicitly add Gaussians for three types of pixels per frame: newly observed, with large color errors, and with large depth errors. We also categorize all Gaussians into stable and unstable ones, where the stable Gaussians are expected to well fit previously observed RGBD images and otherwise unstable. We only optimize the unstable Gaussians and only render the pixels occupied by unstable Gaussians. In this way, both the number of Gaussians to be optimized and pixels to be rendered are largely reduced, and the optimization can be done in real time. We show real-time reconstructions of a variety of large scenes. Compared with the state-of-the-art NeRF-based RGBD SLAM, our system achieves comparable high-quality reconstruction but with around twice the speed and half the memory cost, and shows superior performance in the realism of novel view synthesis and camera tracking accuracy.
@@ 1. Introduction
In this paper, we introduce Real-time Gaussian SLAM (RTG-SLAM), a real-time 3D reconstruction system with an RGBD camera for large-scale environments using Gaussian splatting, featuring a compact Gaussian representation and a highly efficient on-the-fly Gaussian optimization scheme. In our compact Gaussian representation, we force each Gaussian to be either opaque or nearly transparent, with the opaque ones fitting the surface (i.e., depth map) and dominant colors, and transparent ones fitting residual colors. Our intention is to use a single opaque Gaussian to fit a local region of the surface without the need for multiple overlapping Gaussians. However, even for an opaque Gaussian, rendering its depth in the same way as rendering color would produce varying depth values declined from the Gaussian center, making it inaccurate to represent a local area using this Gaussian alone. To this end, we propose to render depth in a different way from color rendering. Following classical point rendering techniques (Zwicker et al., 2001), we treat each opaque Gaussian as an ellipsoid disc on the dominant plane of Gaussian, so that it can well fit a local region or a large flat area by itself. The depth rendering is very convenient under this setting. During color rendering, we already have the sorted Gaussians as well as their opacities for each pixel. By selecting from front to back the first Gaussian whose opacity for the pixel is larger than a given threshold, we consider the ray hits the ellipsoid disc and compute the intersection point using equations of the ray and disc plane. Then the depth for the pixel is equal to the depth of the intersection point. The whole process is differentiable, so Gaussians can be optimized by measuring the differences between the rendered and input depth maps through backpropagation. The compact Gaussian representation can fit the 3D surfaces with much fewer Gaussians, hence largely reducing the memory and computation cost.
@@
[…]
@@ 2. Related Work
There are some concurrent works aiming to integrate 3D Gaussians into dense RGBD SLAM. 3D Gaussians (Kerbl et al., 2023) can render high-quality images in real time, but the optimization is conducted offline typically requiring several minutes. To extend Gaussians to online reconstruction, (Yan et al., 2023) proposes an adaptive expansion strategy to add new or delete noisy 3D Gaussian and a coarse-to-fine technique to select reliable Gaussians for tracking. (Yugay et al., 2023) proposes novel strategies for seeding and optimizing Gaussian splats to extend their use to sequential RGBD inputs. SplaTAM (Keetha et al., 2023) tailors an online reconstruction pipeline to use an underlying Gaussian representation and silhouette-guided optimization via differentiable rendering. (Matsuki et al., 2023) unifies the Gaussian representation for accurate, efficient tracking, mapping, and high-quality rendering. (Huang et al., 2023b) introduces a Gaussian-Pyramid-based training method to progressively learn multi-level features and enhance mapping performance. While promising results are demonstrated, it is still difficult in reaching real-time reconstruction at scale. The reported fastest reconstruction speed is 8.34 fps on an NVIDIA RTX 4090 GPU (Matsuki et al., 2023) on the synthetic Replica dataset (Straub et al., 2019). They did not present the complete reconstruction results on real large scenes either. Thanks to our compact Gaussian representation and highly efficient Gaussian optimization strategy, our method can reconstruct real large scenes in real time with low memory cost.
@@ 4.1. Experimental Setup
We implemented our SLAM system on a desktop computer with an intel i9 13900KF CPU and an Nvidia RTX 4090 GPU. We implemented the mapping and tracking parts in Python using Pytorch framework and wrote custom CUDA kernels for rasterization and back propagation. We used an Azure Kinect as the RGBD camera for real-time scanning. Please refer to the supplementary material for more details.
@@
[…]
@@ 4.2. Evaluation of Online Reconstruction
We can see our reconstruction speed is around twice that of NeRF SLAM methods and about 46×46$\times$46 × that of SplaTAM which is also based on 3D Gaussians. Notably, the memory cost of our method is much smaller compared to other methods, which allows us to scan large-scale environments. Note SplaTAM uses alpha blending to render depths as colors, thus yielding much more Gaussians (7155880 before out of memory in the home scene) than our compact Gaussian representation (987524). Even though they store the RGB values instead of spherical harmonics to reduce the memory overhead, their memory cost is still very high and runs out of memory in the home scene.
@@
We qualitatively compare the rendering quality for novel view synthesis for all methods. The results are shown in Fig. 7. Note the NeRF-based methods require a depth map to synthesize high-quality images, so we use the reconstructed mesh to render depth maps for them. We also quantitatively compare the novel-view synthesis on the ScanNet++ testing views, where the ground truth depth is used for NeRF-based methods, and the results are reported in the supplementary material. We can see that our method and SplaTAM clearly produces higher quality images with much fewer artifacts and higher fidelity appearance. We also quantitatively compare the rendering quality with other methods on the training views of Replica, following Point-SLAM and all concurrent Gaussian SLAM works. Please see the table in the supplementary material. Our method achieves a rendering quality comparable with SplaTAM and Point-SLAM (which needs the ground-truth depth map as input), and consistently outperforms the other NeRF SLAM methods.
@@ Acknowledgements.
The authors would like to thank the reviewers for their insightful comments. This work is supported by NSF China (No. U23A20311 & 62322209), the XPLORER PRIZE, and the 100 Talents Program of Zhejiang University. The source code and data are available at https://gapszju.github.io/RTG-SLAM.
Supplementary Materials
Appendix A Gaussian initialization
Here we introduce how we compute the covariance matrix $\bm{\Sigma}_{\mathbf{u}}$ for each newly added Gaussian for pixel $\mathbf{u}$ in detail. Each new Gaussian $G_{\mathbf{u}}$ is initialized as a flat circle disc. For opaque Gaussians, our goal is to cover the surface as much as possible without largely impacting existing ones. For this, we calculate the distance from the pixel’s 3D position $\mathbf{V}_{k}^{g}(\mathbf{u})$ to its three nearest Gaussians $G_{1,2,3}$ in the scene, and initialize its scale based on the following formula:
$\mathbf{s}_{\mathbf{u},1}=\sqrt{\frac{1}{3}\sum_{i=1}^{3}\left(||\mathbf{V}_{k}^{g}(\mathbf{u})-\mathbf{p}_{i}||-0.5(a_{i}+b_{i})\right)},$
(11) $\mathbf{s}_{\mathbf{u},2}=\mathbf{s}_{\mathbf{u},1}\quad\mathbf{s}_{\mathbf{u},3}=0.1\mathbf{s}_{\mathbf{u},1}.$
Here $a$ is the biggest eigenvalue of the Gaussian covariance matrix and $b$ is the second largest eigenvalue. For the transparent Gaussian, the ratios of its three axes are also set to 1:1:0.1. However, in order to reduce the disruption of the rendering results on other pixels, its maximum scale is limited to 0.01 $m$. Finally, the Gaussian orientation $\mathbf{q}_{\mathbf{u}}$ is initialized such that its shortest axis aligns with the pixel’s global normal $\mathbf{N}_{k}^{g}(\mathbf{u})$. The above initialization method allows the newly added opaque Gaussians to cover the scene surface as much as possible, with little influence on the existing Gaussians in the scene $\mathcal{S}$.
Appendix B Gaussian optimization
Since we only optimize the unstable Gaussians, we use $\mathcal{S}_{unstable}$ to render a light transmission map $\hat{\mathbf{T}}——{unstable}$ before optimization, and the loss $L$ is only computed on the pixels:
(12) $M_{unstable}=\{\mathbf{u}_{unstable}|\hat{\mathbf{T}}_{unstable}(\mathbf{u}_{unstable})<1\}.$
However, in order to achieve fast rendering and optimization, (Kerbl et al., 2023) adopts a tile-based rasterizer for Gaussian splatting: the image is divided into $16\times 16$ tiles, and each Gaussian is assigned a key that combines view space depth and tile ID and then sorted. In fact, when there are just a few pixels within a tile that need to be rendered, instantiating all Gaussians on this tile is quite inefficient. Therefore, we discard those tiles where the number of pixels that need to be rendered is less than $50\%$. This strategy drastically reduces inefficient computation, and increases the overall speed of the optimization process.
Appendix C Implementation Details
We accelerate our system by implementing it in parallel with three threads: one thread for Gaussians optimization, one for front-end online tracking, and the other for back-end graph optimization. The Gaussians optimization and front-end online tracking are implemented by python using the pytorch framework and we write custom CUDA kernels for our rendering process and back propagation. The back-end optimization part is inherited from ORB SLAM2 (Mur-Artal and Tardós, 2017) and implemented in C++. We also build an interactive viewer using the open-source SIBR (Bonopera et al., 2020; Kerbl et al., 2023) to visualize the SLAM process and the reconstructed model. In order to achieve free movement and scanning in the scene, we use a laptop with an intel i7 10750-H CPU and nvidia 2070 GPU connected to an Azure Kinect RGBD camera for data acquisition. The RGBD images captured by the camera are transmitted to the desktop computer through a wireless network and the desktop computer completes the SLAM computations, and then the results are sent back to our viewer on the laptop for visualization. In all our experiments, we sample 5% of pixels for the Gaussians adding. For small-scale synthetic dataset Replica (Straub et al., 2019), we set the optimization time window to 6 and optimize 50 iterations. For our large-scale real Azure dataset, we set the optimization time window to 8 and optimize 50 iterations. For TUM-RGBD (Sturm et al., 2012), we set the optimization time window to 4 and optimize 50 iterations. For ScanNet++ (Yeshwanth et al., 2023), we set the optimization time window to 3 and optimize 75 iterations due to the high-resolution (1752 $\times$ 1168) and sparse viewports. For learning rates, we set $lr_{position}=0.001$, $lr_{SH0}=0.0005$, $lr_{\alpha}=0$, $lr_{scale}=0.004$, $lr_{rotation}=0.001$ on Replica and ScanNet++. And we set $lr_{position}=0.001$, $lr_{SH0}=0.001$, $lr_{\alpha}=0$, $lr_{scale}=0.002$, $lr_{rotation}=0.001$ on our datset and TUM-RGBD. The learning rates of other SH coefficients is 0.05 $\times$ $lr_{SH0}$. For the confidence count threshold, we use $\delta_{\eta}=100$ for synthetic Replica, $\delta_{\eta}=200$ for TUM-RGBD and $\delta_{\eta}=400$ for large-scale Azure and ScanNet++ scenes.
Appendix D Dataset Details
For ScanNet++, we select 4 subsets (8b5caf3398, 39f36da05b, b20a261fdf, f34d532901) for evaluation. The statistics of Azure dataset are listed in Table 4. Note that we don’t have the ground truth for Azure dataset, so it is mainly used for qualitative demonstration (except the evaluation of time & memory performance).
Table 4. Statistics of Azure dataset
Statistic corridor storeroom hotel room home office
Trajectory Length ($m$) 21.9 18.9 39.7 32.2 41.0
Scan Area ($m^{2}$) 43.4 44.3 56.3 69.8 100.2
Frame Number 4890 3310 4838 6130 6889
Appendix E More comparison results
E.1. Quantitative results on different datasets
We add more evaluation on different datasets here. Please note that the low quality 3D model of TUM-RGBD makes it infeasible to evaluate geometry accuracy, as in previous papers. The cameras far apart in ScanNet++ result in tracking failure for classical, NeRF and our SLAM, so only geometry accuracy is evaluated. Our Azure dataset doesn’t have the groundtruth camera or 3D model. The comparison of geometry accuracy on Replica is shown in Table 5. Similar to the paper, our geometry accuracy still outperforms the other methods except Point-SLAM which doesn’t optimize point positions from the perfect depth. SplaTAM and ours degrade in completion because Gaussians cannot complete unscanned regions as NeRF. The comparison of tracking accuracy on Replica are shown in Table 6. Our method consistently outperforms the NeRF SLAM and the concurrent Gaussian SLAM method. We believe this is due to our back-end graph optimization based on ORB landmarks. Please note that in order to ensure fairness in comparison, although there is no noise in the depth images on Replica, we still used frame-to-model ICP, just without applying the bilateral filter to the depth input. We also report the time and memory performance on TUM-RGBD in Table 7. Our system achieves the highest scanning speed and lowest memory cost.
Table 5. Comparison of geometry accuracy on Replica.
[…]
NICE-SLAM(2022b) 2.84 84.44 2.31 84.97
Co-SLAM(2023) 2.33 88.89 1.63 89.94
ESLAM(2023) 1.47 91.44 1.11 93.84
Point-SLAM(2023) 0.61 99.94 2.42 86.85
SplaTAM(2023) 2.88 73.89 3.57 71.68
Ours 0.75 98.87 2.81 82.76
Table 6. Comparison of tracking accuracy (unit: $cm$) on Replica.
Method Rm 0 Rm 1 Rm 2 Off 0 Off 1 Off 2 Off 3 Off 4 Avg.
NICE-SLAM(2022b) 0.97 1.31 1.07 0.88 1.00 1.06 1.10 1.13 1.06
Co-SLAM(2023) 0.69 0.59 0.73 0.87 0.47 2.16 1.30 0.62 0.93
ESLAM(2023) 0.66 0.62 0.55 0.44 0.43 0.50 0.66 0.53 0.55
Point-SLAM(2023) 0.54 0.43 0.34 0.36 0.45 0.44 0.63 0.72 0.49
SplaTAM(2023) 0.47 0.42 0.32 0.46 0.24 0.28 0.39 0.56 0.39
Ours 0.20 0.18 0.13 0.22 0.12 0.22 0.20 0.19 0.18
Finally we compare the rendering quality. The results of training view synthesis quality on Replica are reported in Table 8. Please note that this comparison is actually unfair as Point-SLAM (2023) takes the ground-truth depth maps as input to help sampling the 3D volume for rendering. In contrast, our method and SplaTAM (2023) do not require any auxiliary input. Even so, our method still achieves a rendering quality comparable with Point-SLAM and SplaTAM, and consistently outperforms the other NeRF SLAM methods. The quantitative comparison of novel-view synthesis on ScanNet++ testing views is reported in Table 9. Our method is comparable to SplaTAM and outperforms the other NeRF-SLAM methods.
Table 7. Comparison of time and memory performance on TUM-RGBD.
[…]
NICE-SLAM(2022b) 0.06 9930
CO-SLAM(2023) 7.18 18607
ESLAM(2023) 0.30 18617
Point-SLAM(2023) 0.26 11000
SplaTAM(2023) 0.14 12100
Ours 21.74 3563
Table 8. Comparison of train view synthesis on Replica.
Method Metric Rm 0 Rm 1 Rm 2 Off 0 Off 1 Off 2 Off 3 Off 4 Avg.
[…]
Figure 9. Comparison of novel view synthesis on Replica.
Table 9. Comparison of novel view synthesis on ScanNet++.
[…]
NICE-SLAM(2022b) 23.71 0.797 0.341
CO-SLAM(2023) 23.20 0.837 0.413
ESLAM(2023) 27.06 0.856 0.322
Point-SLAM(2023) 21.85 0.802 0.404
SplaTAM(2023) 27.77 0.864 0.233
Ours 27.27 0.872 0.295
E.2. Comparison with more SLAM systems
We show more comparisons with ElasticFusion (Whelan et al., 2015), BundleFusion (Dai et al., 2017), and GO-SLAM (Zhang et al., 2023). The tracking accuracy is evaluated on Replica and TUM-RGBD and the results are listed in Table 10. we also evaluate the geometric accuracy on Replica as shown in Table 11. Our method achieves comparable geometry accuracy as BundleFusion. GO-SLAM’s completion numbers are worse than those of its paper due to considering unscanned regions for fair comparison.
Table 10. Comparison of tracking accuracy (unit: $cm$) with more SLAM systems.
Method Replica TUM-RGBD
ElasticFusion(2015) 1.13 2.07
BundleFusion(2017) 0.46 1.63
GO-SLAM(2023) 0.37 1.28
Ours 0.18 1.06
Table 11. Comparison of geometry accuracy on Replica with more SLAM systems.
[…]
ElasticFusion(2015) 1.13 96.33 4.43 75.25
BundleFusion(2017) 0.77 99.88 5.35 76.69
GO-SLAM(2023) 2.51 76.93 5.11 65.10
Ours 0.75 98.87 2.81 82.76
Appendix F More ablation studies
F.1. Ablation study on sampled pixel number
Here we evaluate the influence of the number of sampled pixels for adding Gaussians. We set the sampling ratio to 5%, 10%, and 20% for each frame to reconstruct Replica office0 and reported the image quality metrics. The results suggest that even if we sample a small number of images for reconstruction, the image quality is not significantly affected.
Table 12. Ablation study on sampled pixel number.
[…]
$5\%$ 39.01 0.965 0.072
$10\%$ 39.63 0.970 0.051
$20\%$ 39.31 0.972 0.044
F.2. Ablation study on stable/unstable Gaussians
We test the impact of our proposed stable/unstable Gaussians on time performance. We report the optimization time per iteration, for optimizing all Gaussians using the whole image, optimizing only unstable Gaussians using the whole image, and optimizing only unstable Gaussians using only the pixels covered by them. As seen in Table 13, our strategy greatly improves the optimization speed.
Table 13. Ablation study on stable/unstable Gaussians.
Scene Storeroom Hotel room Home
$\mathcal{S}$ 9.8ms 8.4ms 8.9ms
$\mathcal{S}_{unstable}$, all pixels 7.4ms 6.5ms 6.4ms
$\mathcal{S}_{unstable}$, unstable pixels 5.2ms 4.7ms 4.3ms
F.3. Ablation study on depth rendering
Our depth blending is tightly coupled with our Gaussian adding and state management, so in the paper we show that alpha blending yields much more Gaussians through the comparison with SplaTAM which uses alpha blending (987524 vs 7155880). Here for better ablation study, we first use our depth blending to determine the adding of opaque/transparent Gaussians as well as the states, and then replace our depth blending with alpha blending for optimization. The results are listed in Table 14. We can see with similar Gaussian numbers, our depth blending outperforms alpha blending in terms of geometry accuracy and tracking accuracy.
Table 14. Ablation study on depth rendering.
[…]
Alpha-blending 2.48 70.54 3.32 75.01 1.24 468916
Ours 0.75 98.87 2.81 82.76 0.18 431692
F.4. Ablation study on confidence count threshold
The ablation study on confidence count threshold $\delta_{\eta}$ on Replica is shown in Table 15. As $\delta_{\eta}$ increases, the Gaussians will be in the unstable state for a longer time, resulting in a slower speed. On the other hand, if $\delta_{\eta}$ is small, the Gaussians may not be fully optimized, yielding reduced rendering quality.
Table 15. Ablation study on confidence count threshold.
[…]
50 33.63 18.21
100 35.43 17.31
200 35.12 16.59
400 35.37 15.49
F.5. Ablation study on backend pose optimization
Here we evaluate the effect of backend pose optimization. We test the tracking accuracy on Replica and TUM-RGBD, and report the results in Table 16. We also test the geometry accuracy on Replica and report the results in Table 17. On high-quality images on Replica, we achieve relatively high accuracy using only the frontend ICP. However, on low-quality images on TUM-RGBD, we rely more on the ORB backend, because the ICP may be performed on the Gaussians still under optimization.
Table 16. Ablation study on backend pose optimization in terms of tracking accuracy .
Method Replica TUM-RGBD fr1_desk TUM-RGBD fr2_xyz
ElasticFusion without backend 0.68 2.93 1.32
ElasticFusion 1.13 2.53 1.17
Ours without backend 0.22 5.39 1.63
Ours 0.18 1.66 0.38
Table 17. Ablation study on backend pose optimization in terms of geometry accuracy .
[…]
ElasticFusion without backend 1.05 98.32 4.33 77.69
ElasticFusion 1.13 96.33 4.43 75.25
Ours without backend 0.95 97.57 2.74 83.17
Ours 0.75 98.87 2.81 82.76
F.6. Ablation study on SH coefficients
Here we report the rendering quality using SHs and RGB colors. We report the results of training view synthesis on Replica in Table 18. We also report the results of novel view synthesis on ScanNet++ in Table 19. We can notice that using SH coefficients has better rendering quality.
Table 18. Ablation study on SH coefficients in terms of training view synthesis.
[…]
RGB 33.90 0.951 0.113
SH 35.43 0.982 0.109
Table 19. Ablation study on SH coefficients in terms of novel view synthesis .
[…]
RGB 25.67 0.855 0.304
SH 27.27 0.874 0.291
F.7. Ablation study on outlier pruning
Here we evaluate the influence of our outlier pruning strategy. We report the number of Gaussians every 1000 frames in the Azure hotel room scene in Table 20 and the results show that the Gaussian number will increase significantly without outlier pruning.
Table 20. Ablation study on outlier pruning.
Frame ID Ours Ours without outlier pruning
1000# 151947 211390
2000# 268625 382736
3000# 507987 797242
4000# 675818 1073366